OpenStudy (anonymous):

The graph of the following system of equations is -2x + y = 3 4x + 2y = 2 Answer Overlapping lines Parallel lines Intersecting lines

OpenStudy (anonymous):

Happy thanksgiving ya'll.

hero (hero):

y = 3 + 2x y = 1-2x Graph them and see

OpenStudy (anonymous):

i don't know how to graph them...

OpenStudy (ash2326):

intersecting lines

OpenStudy (muhammad_nauman_umair):

intersecting

OpenStudy (muhammad_nauman_umair):

As, line#1 -2x+y=3 \[for x=0;y=3 fory=0;x=-1.5points for line#1 are (0,3),(-1.5,0)\] for line#2 4x+2y=2 for x=0;y=1 fory=0;x=0.5 points for line#2 are (0,1),(0.5,0) plot two lines on the graph and u will c it\

OpenStudy (muhammad_nauman_umair):



OpenStudy (muhammad_nauman_umair):

Got it or not?...

hero (hero):

She just wants the answers man. She doesn't care